OpenStudy (anonymous):

What is the simplest form of the radical expression? (sqrt(2)+sqrt(5))/(sqrt(2)-sqrt(5)

OpenStudy (anonymous):

\[\sqrt{2}+\sqrt{5}\div \sqrt{2}-\sqrt{5}\]